OpenStudy (anonymous):

what is the amplitude, period, and phase shift of f(x)=-3cos(4x+pie)+6 ? i need help step by step please

OpenStudy (anonymous):

@perl

OpenStudy (anonymous):

@dtan5457

OpenStudy (anonymous):

Let's consider a more general case. Take \(f(x)=a\cos (b(x-c))+d\). The amplitude is \(|a|\), the period is \(\dfrac{2\pi}{b}\), the horizontal/phase shift if \(c\), and the vertical shift is \(d\). (The same works for \(\sin\)).

OpenStudy (anonymous):

i dont quite understand...((omg your a sith lol starwars is awesome))

OpenStudy (anonymous):

@SithsAndGiggles

OpenStudy (anonymous):

If we want to make a direct connection between the general form and the particular function you're given, it's a matter of writing the function accordingly. It should be clear to you that \(a=-3\). What's the value of \(|a|\)? (This is your amplitude.) To find the period, we need the coefficient of \(x\) to be \(1\), but we're given \(4\). Factorizing, we get \[4x+\pi=4\left(x+\frac{\pi}{4}\right)\] so we have that \(b=4\), and additionally that \(c=-\dfrac{\pi}{4}\). What are the period and phase shifts going to be? You don't have to worry about the vertical shift \(d\), but it's apparent that \(d=6\) and so the curve is shifted \(6\) units up.

OpenStudy (anonymous):

so would my answer be amplitude = 3 period= pi/2 phase shift : x = - pi/4

OpenStudy (anonymous):

@sithsandgiggles

OpenStudy (anonymous):

@dtan5457

OpenStudy (anonymous):

@GracieBugg

OpenStudy (graciebugg):

There is no way that I can answer this ,I suck that bad at math xD sorry!

OpenStudy (anonymous):

I'm sorry @hailhydra I really wish I could help, but I haven't learnt that, yet. I'm soo sorry

OpenStudy (anonymous):

@dan815

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (anonymous):

aww , thanks for letting me know though haha :) @Music_me

OpenStudy (anonymous):

Yes your answer is right.

OpenStudy (anonymous):

thank you so much @SithsAndGiggles so so so much!~

OpenStudy (anonymous):

@sithandgiggles

OpenStudy (anonymous):

*sithsandgiggles

OpenStudy (anonymous):

@sithsandgiggles

OpenStudy (anonymous):

wait! i just noticed isnt the amplitude equal -3 ? i thinky ou misread it

OpenStudy (anonymous):

@dan815

OpenStudy (anonymous):

@dan815

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

@amistre64

OpenStudy (anonymous):

@aaronq

OpenStudy (anonymous):

@jim_thompson5910 can you help clarify this last part for me?

jimthompson5910 (jim_thompson5910):

what's your question?

OpenStudy (anonymous):

what is the amplitude, period, and phase shift of f(x)=-3cos(4x+pie)+6 ?

jimthompson5910 (jim_thompson5910):

how much do you have so far

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

hint: y = A*cos(Bx - C) + D |A| is the amplitude T = 2pi/B is the period C/B is the phase shift y = D is the midline

jimthompson5910 (jim_thompson5910):

oh I see you wrote `so would my answer be` ` amplitude = 3 ` ` period= pi/2` ` phase shift : x = - pi/4`

jimthompson5910 (jim_thompson5910):

amplitude = 3 is correct period= pi/2 is correct For the phase shift, I would simply just say "-pi/4 units"

OpenStudy (anonymous):

so would my answer be amplitude = -3 period= pi/2 phase shift : x = - pi/4

OpenStudy (anonymous):

isnt the amplitude -3?

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

no the amplitude is always positive it's a distance

jimthompson5910 (jim_thompson5910):

|A| = |-3| = 3

OpenStudy (anonymous):

ohhhhhhhhhhhhhhhhh okk thank you i checked it in my notes thank you @jim_thompson5910

jimthompson5910 (jim_thompson5910):

np

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

@amistre64

OpenStudy (anonymous):

hey, i was wondering if you can explain to me how the period is found?

OpenStudy (amistre64):

well, what is your concept of periodic?

OpenStudy (anonymous):

like it measures time?

OpenStudy (amistre64):

1 period is defined as the time it takes to start over again ..... agreed?

OpenStudy (amistre64):

if we take the smallest interval as the fundamental periodic rate then... we have to ask outselves; when does trig(0) = trig(a) ?? then the period is defined as 'a'

OpenStudy (amistre64):

you have cos(u) for whatever the inside is ... what is the first time that cos(0) = cos(a) ??

OpenStudy (anonymous):

im not grasping the understanding of your question :(

OpenStudy (amistre64):

wht is the value of cos(0) ?

OpenStudy (anonymous):

is it 0 or 1?

OpenStudy (anonymous):

could it be that there is no cos ?

OpenStudy (amistre64):

no there is a cos, for whatever that means and its not my responsibility to know cos(0) at the moment ... you tell me the value of cos(0) and ill correct you.

OpenStudy (anonymous):

is it equal to 1?

OpenStudy (amistre64):

cos(0) = 1 now, the period of cos is the value 'a' such that cos(a) = 1 for the first time again .... assume x is on an interval from 0 to a

OpenStudy (amistre64):

is cos(pi/4) = 1 ? is cos(pi/3) = 1 ? is cos(pi/2) = 1 ? is cos(pi) = 1 ? is cos(3pi/2) = 1 ? is cos(2pi) = 1 ?

OpenStudy (anonymous):

is it is cos(pi/2) = 1

OpenStudy (amistre64):

good, then cos(0) = cos(2pi) is the smallest interval in which cos repeats itself. cos(u) has a normal period of 2pi now, lets use this to determine the distance in your problem let u = 4x+pi when u1 = 0 when x=? when u2 = 2pi when x=? and therefore the period is u2-u1

OpenStudy (anonymous):

isnt the distance the amplitude = 3

OpenStudy (amistre64):

of course not, the distance along the x axis does not measure height (amplitude)

OpenStudy (anonymous):

im a little confused

OpenStudy (amistre64):